A monkey climbs and another monkey climbs down a rope hanging from a tree with same uniform acceleration separately.If the respective masses of monkeys are in the ratio 2:3 the common acceleration must be

Grade:11

2 Answers

Arun
6 years ago

Gravity is the force by which earth attracts all bodies towards itself. It is same for all bodies and does not depend upon the mass of body.

hence common acceleartion must be g.

Tiya
4 years ago
g/5
a=(m2-m1)g/(m2+m1)
a=(3-2)g/3+2
a=g/5
This formula is given as per pulley problem.